Al quit his job as a manager at a Savings bank where he earned $90,000 a year. He cashed in $300,000 in corporate bonds that earned 1% interest annually to set up a gym. Al has decided to buy a store front and setup exercise classes. There are 900 people who will pay $1250 a year for unlimited classes; $980 from each person goes for instructors, rent, maintenance, equipment, insurance, depreciation etc.

Write your answers to the following questions on the answer sheet.

A. What are Al’s total revenues? (3 pts)

B. What are Al’s explicit costs? In numbers (3 pts)

C. What is his accounting profit? Numbers (3 pts)

D. List 2 in numbers 2 implicit costs that Jon has not included. (3 pts)

E. What is Al ’s pure economic profit (or loss) in numbers? (3 pts
Answer #1

A) Total revenue = 900*1250 = 1125000

B) Explicit costs = 980*900 = 882000

C) Accounting profit = Total revenue - explicit costs = 1125000-882000 = 243000

D) Implicit costs = salary foregone+interest on bonds

= 90000+300000*1%

= 93000

E) Economic profit = Accounting profit - Implicit costs

= 243000-93000

= 150000